Suppose that 15% of all Americans use CNN as their primary source of news. A random sample of 100 Americans is selected. What is the probability that fewer than 22 people in your sample of 100 Americans will use CNN as their primary news source?
Answer #1

solution:-

given that p = 0.15 , n = 100

mean = n*p = 100*0.15 = 15

standard deviation = sqrt(n*p*(1-p)) = sqrt(100*0.15*(1-0.15)) = 3.5707

here asked that

=> P(x < 22)

=> P(z < (22-15)/3.5707)

=> P(z < 1.96)

=> 0.9750